OpenStudy (anonymous):

The position of an object time t is given by s(t)=7-14t . Find the instantaneous velocity at t=7 by finding the derivative. Please explain!

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

The derivative is the slope of the function. You can plug any x value into the derivative, and get the slope of the original equation at that point. Note: Instantaneous .... is a fancy way of saying find the slope! Apply the chain rule since we have a linear equation. s'(t) = -14 By the chain rule 7 goes to 0. -14t^1 becomes 1 * -14t^(1 -1) = -14t^0 = -14. This means that the instantaneous velocity at any point is the same.

OpenStudy (anonymous):

So for example, in a problem like s (t)=-9-5t with t=4. I would use the chain rule with -5?

OpenStudy (anonymous):

@Iamsinged